OpenStudy (anonymous):

Solve the following system by the comparison method, x-4y-1=0, x+5y-4=0

OpenStudy (anonymous):

1 + 4y = 4 - 5y, I think.

OpenStudy (anonymous):

3] x = 1 + 4y x = 4 - 5y so 1 + 4y = 4 - 5y..........4y + 1 = 4 - 5y

OpenStudy (anonymous):

i think